Hama is three times As old as David and the sum of their ages is 52. Determine their ages
Answer by ikleyn(52803)

.

D + 3D = 52

4D = 52

D = 52/4 = 13 for David.     ANSWER

3*13 = 39 for Hama.          ANSWER